The weight X of a lion in the wild has a normal distribution with mean 345 pounds and a standard deviation of 75 pounds. What proportion (i.e., probability) of the lion is between 200 and 500 pounds? Use the Z-score table!
Answer #1

Solution :

Given that ,

mean = = 345

standard deviation = = 75

P(200 < x < 500) = P((200 - 345)/ 75) < (x - ) /  < (500 - 345) / 75) )

= P(-1.93 < z < 2.07)

= P(z < 2.07) - P(z < -1.93)

= 0.9808 - 0.0268 Using standard normal table,  

Probability = 0.9540

The probability of the lion is between 200 and 500 pounds is 0.9540.